Harmonic analysis methods for incompressible flows of an ideal fluid

Mikhail Vishik The University of Texas at Austin

We formulate uniqueness results for the Euler equations of an ideal incompressible fluid and give some examples of (classes of) flows in terms of their Littlewood-Paley decomposition.